# B2N3D: Progressive Learning from Binary to N-ary Relationships for 3D Object Grounding

### v1

@@ Abstract
Localizing 3D objects using natural language is essential for robotic scene understanding. The descriptions often involve multiple spatial relationships to distinguish similar objects, making 3D-language alignment difficult. Current methods only model relationships for pairwise objects, ignoring the global perceptual significance of n-ary combinations in multi-modal relational understanding. To address this, we propose a novel progressive relational learning framework for 3D object grounding. We extend relational learning from binary to n-ary to identify visual relations that match the referential description globally. Given the absence of specific annotations for referred objects in the training data, we design a grouped supervision loss to facilitate n-ary relational learning. In the scene graph created with n-ary relationships, we use a multi-modal network with hybrid attention mechanisms to further localize the target within the n-ary combinations. Experiments and ablation studies on the ReferIt3D and ScanRefer benchmarks demonstrate that our method outperforms the state-of-the-art, and proves the advantages of the n-ary relational perception in 3D localization.
Figure 1: The blue part illustrates a reasoning process for 3D object grounding with multiple referential relationships. The gray part shows the main difference between the previous methods and ours. Compared to the previous methods, which only perceives the relationships of paired objects, our method extends relationship modeling from binary to n-ary, reducing redundant or erroneous relationship information during multi-modal fusion.
@@ I Introduction
Anumber of real-world robotic tasks require natural language instructions to propose actions in the real world [1]. Since 3D representations can provide richer information, research in 3D vision-language is advantageous for robotic applications [2]. 3D object grounding is a crucial task in the 3D vision-language field, aimed at localizing an object in a 3D scene based on textual descriptions. [3, 4].
@@
A real-world scene often contains multiple similar objects and has a complex spatial arrangement, people must describe the specific target through referential relationships. As shown in Figure 1, understanding the spatial relationships between “plant-bookcase” and “plant-desk” is necessary for localizing the described target among the five “plant”. Due to the inherent differences between the visual and textual modalities [5, 6], simultaneously mapping multiple relationships in the text to entities is a significant challenge for 3D object grounding. Recent methods mainly use Transformer-based frameworks to achieve 3D-language alignment for localization [7, 8, 9, 10, 11, 12, 13, 14]. Some works [15, 16, 17, 18, 19] create scene graphs with entities as nodes and aggregate information from neighboring or text-related nodes to learn entity relationships implicitly. Others [13, 20, 10] integrate independent geometric modules into transformers to enhance spatial relational representation. However, when using multiple referential relationships to differentiate the target from similar distractors, current methods still often fail to localize the correct object due to limited relational perception.
@@
Many descriptions, such as the example in Figure 1, involve multiple entities and relationships. Most existing methods perform relational perception only from pair-wise entities ((a) in Figure 1). The multi-modal fusion utilizes all relational information that corresponds with the local description (both “plant_9 and desk_44” and “plant_62 and desk_43” fulfill “the plant tucked in the furthest corner of the desk”). This may increase the uncertainty of the object representation. In cases where multiple entity pairs simultaneously adhere to the described relationships, local reasoning is insufficient for precise localization.
@@
Although graph-based models can represent both local neighborhoods and global contexts [21], the scene graphs in current research establish connections between all spatially adjacent or semantically related entities. This also produces incorrect guidance in cross-modal understanding due to the aggregation of excessive noisy information in graph nodes. Our goal is to infer the n-ary combinations that simultaneously satisfy all referential conditions detailed in the description. This process involves a close relationship between the multi-modal fusion feature and n-ary global representation, while ensuring the removal of interference from noisy information.
@@ II Related Work
The research on 3D object grounding aims to localize the specific object using referential descriptions in 3D scenes. The current mainstream methods rely on the two benchmarks ScanRefer [3] and ReferIt3D [4] for further research. From a modeling structure perspective, these methods can be classified as two-stage and one-stage [22]. The two-stage method first employs a 3D detector to identify object proposals and then utilizes a vision-language perception model to determine the target [13, 11, 23]. The one-stage method processes scene point clouds end-to-end and directly outputs the target position [24, 25, 20]. Describing complex scenes often requires multiple spatial relationships to clarify object positions, making the accurate understanding of these relationships a primary issue. Zhao et al. [10] and Chang et al. [13] integrate pre-extracted spatial relationship matrices into the transformer to improve relation awareness. Cai et al. [20] improve geometric relationship expressions through the proposed attribute-relation enhancement module. Huang et al. [18] construct an instance graph to aggregate neighboring relationship information. Chen et al. [16] point out that modeling only neighboring objects in the graph ignores the distance expression of relationships and propose a spatial self-attention layer. The state-of-the-art methods [14, 12] focus more on better visual representations of objects from attributes or semantic labels. Additionally, some works utilize large language models for text analysis or data augmentation to enhance the language comprehension capabilities of models [26, 27, 28, 29]. However, these methods still struggle to make accurate distinctions when faced with multiple objects identical to the target category. In this paper, we explicitly model binary and n-ary relations to facilitate global multi-modal alignment, which enables more accurate localization of objects under complex circumstances.
@@ II-B Scene Graph in Vision-Language
Scene graphs provide a high-level understanding of visual scenes, typically representing objects as nodes and relational representations as edges [21]. Because of the topological homogeneity of visual and textual modalities in scene graphs, scene graphs are widely employed in vision-language tasks to facilitate cross-modal alignment and joint understanding [30, 31, 32]. Graph neural networks (GNNs) can automatically learn the relationships between nodes when paired with scene graph structures. The introduction of the graph attention network (GAT) [33], which incorporates attention attributes, helps understand complex scenes [34, 35, 36, 37]. In the field of 3D visual understanding, researchers also construct scene graphs to enhance model understanding of geometric relationships [38, 39, 40]. Wei et al. [41] generate 3D scene graphs directly from raw point clouds in an end-to-end manner. Koch et al. [42] use the objects detected by the 3D detector as graph nodes and construct scene graphs with LLM filtering object-pair relationships as edges. Chen et al. [43] predict the relationship score by concatenating the features of two objects and inputting them into a fully connected network, which is used to construct edges. Strader et al. [44] utilize logical rules in spatial ontology as prior knowledge, employing LLM to extract common spatial concepts and establish a scene graph. However, the above methods only focus on the spatial or semantic relationships between pairwise objects when constructing scene graphs. They cannot effectively adapt to the 3D localization task with multiple referential relationships. Our method predicts n-ary relationships globally to establish graph edges and enhance vision-language alignment.
@@ III-A Overview
The overall structure of our method is shown in Figure 2. Descriptions and objects within 3D scenes are encoded into text and object features, respectively. Subsequently, the binary-to-n-ary progressive relational learning module (B2N-PRL) module anticipates entity relationships via a text-guided binary to n-ary modelling process. During the training phase, we employ the LLM to extract entity combinations from descriptions related to the referential dependencies, which are used to supervise the progressive relational learning. The top $K_{2}$ n-ary combinations are used to construct scene graphs. The attention-driven graph learning module updates node features and localizes the target by confidence from a multilayer perceptron (MLP).
@@ III-B Object-Level Representation
Since entities are the basic units of relational learning, we prioritize object-level visual feature encoding within scene contexts. Following the two-stage detect-then-match approaches, the scene point cloud is decoded into instance-level point clouds via segmentation networks, then extracts object features using a 3D point cloud encoder. Inspired by [19], our method employs a training-free segmentation model based on multiple views to generate class-agnostic object proposals. Simultaneously, we harness projection data during the segmentation process to derive 2D features from the pre-trained multi-modal model. The incorporation of pre-trained 2D visual features addresses the inadequate representation of surface color and texture details in point cloud encoding. This enables the model to effectively assimilate comprehensive multi-modal attributes within the constraints of limited 3D data. Therefore, we fuse the 2D features with 3D object features to enhance object-level visual representation as follows:
@@
where $F_{obj\_2D}$ and $F_{obj\_3D}$ represent the 2D object features from the pre-trained model CLIP [45] and the 3D object features from the point cloud encoder, respectively. $F_{obj\_2D}$ is mapped by linear function $\phi$ to obtain fine-tuned features that have the same dimensions as $F_{obj\_3D}$. The MLP-based fusion of point cloud features with multi-modal 2D features may potentially obscure certain geometric information. To mitigate this issue, the enhanced features are subsequently integrated with the original point cloud features.
@@ III-C Progressive Relational Learning
The textual descriptions for the target with distractors often involve multiple referential relationships, which are also seen as the combination of several pairwise object relationships. However, existing models predominantly employ pairwise relationship modeling followed by direct fusion with text modalities, exhibiting insufficient global contextual awareness in descriptions with n-ary relational dependencies. We reformulate the object grounding task as a multi-relational structure inference problem. The target entity is localized by identifying candidate object combinations that satisfy the composite relational constraints. We utilize the LLM to extract binary relationships with textual representation about entity names. The corresponding objects in the scene are identified using these entity names to supervise the relational learning model. We design the progressive relational learning module, B2N-PRL, for reasoning about binary to n-ary relationships, as shown in Figure 3.
@@ III-C2 Binary Relational Modeling
Binary relations as fundamental units for modeling complex spatial relationships are encoded from paired objects. We employ a BERT-based [46] textual encoder to process natural language sentences and obtain text features $T$. During the perception process, sentence decomposition is unnecessary, and the holistic textual representation T directly engages in multi-modal interactions with visual information at various stages. Similarly to previous work [13], we improve the spatial perception ability of the model by integrating box embedding $F_{box}$ and artificially extracted geometric features $F_{geo}$. The specific calculation is as follows:
$\mathbf{O^{\prime}}_{i}=\mathbf{Att}_{cross}(\mathbf{O}_{i}+F_{box-i},\mathbf{T})$ (2)
$\mathbf{B}_{i,j}=\mathbf{O^{\prime}}_{i}\times\mathbf{O^{\prime}}_{j}+F_{geo-ij}$ (3)
@@
$L_{br}=-\frac{1}{N_{1}}\sum_{i=1}^{N_{1}}\left[r_{i}\log(S_{1,i})+(1-r_{i})\log(1-(S_{1,i})\right]$ (4)
where $N_{1}=K_{1}\times K_{1}$ represents the number of binary combinations, and $r_{i}\in\{0,1\}$ indicates whether the $i$-th entity combination matches the description.
@@ III-D2 Graph-Structured Multi-Modal Network
We integrate 3D vision with text in a graph structure and propose a multi-modal attention network. As shown in Figure 4, we employ cross-attention to facilitate interaction between object features and text features, and graph attention to aggregate node information concerning referential relationships through n-ary connections. Specifically, node features are fused with binary features obtained from the relation learning module using a self-attention mechanism to capture object-to-object dependencies. Then we utilize the multi-head cross-attention operation to fuse the textual information and achieve global perception within a graph attention layer. The graph attention operation updates the node values as follows:
@@ III-E Loss Function
During training, we set up the localization loss $L_{ref}$ to supervise the final target prediction results and use the classification loss $L_{t}$ and $L_{v}$ to guide the textual and visual encoding as in previous work [30, 26, 13, 14]. It is noted that previous methods use original instance classes as the gt labels for $L_{v}$. We reduce the number of mapped classes to rely less on the classification information inherent in the visual modality. In addition to them, we propose binary and n-ary relational losses ($L_{br}$ and $L_{nr}$ in Section III-C) to supervise the B2N-PRL module. The whole loss function is grouped as:
@@ IV-B Implementation Details
We use GPT-4o mini [47] as the LLM-based agent to extract entity combinations related to spatial relationships from descriptions. All multi-head cross-attention, self-attention, and graph attention layers in our methods are configured with eight heads. The loss weight coefficients $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ are set to 0.1, 0.5, and 2.0, respectively. The models are trained with a batch size of 20 for 150 epochs, using the Adam [48] optimizer. The learning rate is initialized to $5\times 10^{-4}$ with a decay of 0.65 every 10 epochs from 30 to 80. In the B2N-PRL module, the number of selected binary and n-ary relationships is set to $K_{1}$=16 and $K_{2}$=16.
@@
ReferIt3D [4] 2020 35.6 27.9 43.6 32.5 37.1 40.8 31.5 44.7 39.2 40.8
TGNN [18] 2021 37.3 30.6 44.2 35.8 38.0 45.0 36.9 48.5 45.8 45.0
3DVG-Trans [10] 2021 40.8 34.8 48.5 34.8 43.7 51.4 44.9 54.2 44.6 51.7
FFL-3DOG [17] 2021 41.7 35.0 48.2 37.1 44.7 - - - - -
LAR [49] 2022 48.9 42.3 58.4 47.4 52.1 59.4 51.2 63.0 50.0 59.1
SAT [9] 2021 49.2 42.4 56.3 46.9 50.4 57.9 50.0 61.2 49.2 58.3
M3DRef-CLIP [23] 2023 49.4 43.4 55.6 42.3 52.9 - - - - -
3D-SPS [25] 2022 51.5 45.1 58.1 48.0 53.2 62.6 65.4 56.2 49.2 63.2
EDA [24] 2023 52.1 - - - - 68.1 - - - -
BUTD-DETR [50] 2022 54.6 48.4 60.7 46.0 58.0 67.0 63.2 68.6 53.0 67.6
MVT [11] 2022 55.1 49.1 61.3 54.3 55.4 64.5 58.8 66.9 58.4 64.7
ViewRefer [26] 2023 56.0 49.7 63.0 55.1 56.8 67.0 62.1 68.9 52.2 67.7
ViL3DRel [16] 2022 64.4 57.4 70.2 62.0 64.5 72.8 67.9 74.9 63.8 73.2
MiKASA [13] 2024 64.4 59.4 69.7 65.4 64.0 75.2 67.3 78.6 70.4 75.4
MA2TransVG [12] 2024 65.2 57.6 71.1 62.5 65.4 73.9 69.3 76.0 64.5 73.8
xM_Match [14] 2024 66.2 59.9 72.8 63.8 67.5 74.6 71.3 75.9 65.0 74.7
@@
3D-SPS [25] one-stage Pixel-N 29.8 37.0
BUTD-DETR [50] one-stage - 32.8 37.1
3DVLP [51] one-stage Pixel-N 33.4 40.5
EDA [24] one-stage - 37.6 42.3
MCLN [52] one-stage - 38.4 42.6
TGNN [18] two-stage 23.2 29.7
InstanceRefer [15] two-stage - 24.8 32.9
ViewRefer [26] two-stage - 26.5 33.7
ViL3DRel [16] two-stage - 30.7 37.7
3D-LP [53] two-stage 32.2 38.1
xM_Match [14] two-stage - - 39.3
@@
ScanRefer [3] two-stage Pixel-N 21.1 27.4
3DVG-Trans [10] two-stage Pixel-N 28.4 34.7
3DJCG [20] two-stage Pixel-N 30.8 37.3
D3Net [54] two-stage Pixel-N 30.1 37.9
3D-LP [53] two-stage Pixel-N 33.4 39.5
MA2TransVG [12] two-stage Pixel-N 41.4 45.7
M3DRef-CLIP [23] two-stage Object-3 36.8 44.7

### v2

@@ Abstract
Localizing the object with natural language descriptions is essential in 3D scene understanding. The descriptions often involve multiple spatial relationships to distinguish similar objects, making 3D-language alignment difficult. Current methods only model relationships for pairwise objects, ignoring the global perceptual significance of n-ary combinations in multi-modal relational understanding. To address this, we propose a novel progressive relational learning framework for 3D object grounding. We extend relational learning from binary to n-ary to identify visual relations that match the referential description globally. Given the absence of specific annotations for referred objects in the training data, we design a grouped supervision loss to facilitate n-ary relational learning. In the scene graph created with n-ary relationships, we use a multi-modal network with hybrid attention mechanisms to further localize the target within the n-ary combinations. Experiments and ablation studies on the ReferIt3D and ScanRefer benchmarks demonstrate that our method outperforms the state-of-the-art, and proves the advantages of the n-ary relational perception in 3D localization.
@@ I Introduction
Inrecent years, applying natural language to 3D scene understanding has become a hot topic [yan2023comprehensive]. Tasks involving reasoning about real-world scenes, such as 3D dense captioning [mao2023complete] and visual question answering [ye20223d], have garnered significant attention. 3D object grounding aims at associating textual descriptions with specific objects [chen2020scanrefer, achlioptas2020referit3d]. It requires building a vision-language model to reason about object contextual relationships.
Figure 1: The target or referred objects mentioned in the description often have multiple similar instances present in the scene like “plant” and “desk” in the figure. Accurately localizing the target needs to identify the correct combination of objects that jointly satisfies the multi-relational constraints (“plant_9 desk_43 bookcase_34”). Compared to the previous methods, which only perceives the relationships of paired objects, our method employs a progressive approach to learn relational composition, achieving global perception of n-ary relationships.
@@
A real-world scene often contains multiple similar objects and has a complex spatial arrangement, people must describe the specific target through referential relationships. As shown in Fig. 1, understanding the spatial relationships between “plant-bookcase” and “plant-desk” is necessary for localizing the described target among the five “plant”. Due to the inherent differences between the visual and textual modalities, simultaneously mapping multiple relationships in the text to entities is a significant challenge for 3D object grounding. Recent methods mainly use Transformer-based frameworks to achieve 3D-language alignment for localization [he2021transrefer3d, roh2022languagerefer, yang2021sat, zhao20213dvg, huang2022multi, xu2024multi, chang2024mikasa, zhang2024cross]. Some works [yuan2021instancerefer, chen2022language, feng2021free, huang2021text, xiao2025as3d] create scene graphs with entities as nodes and aggregate information from neighboring or text-related nodes to learn entity relationships implicitly. Others [chang2024mikasa, cai20223djcg, zhao20213dvg] integrate independent geometric modules into transformers to enhance spatial relational representation. However, when using multiple referential relationships to differentiate the target from similar distractors, current methods still often fail to localize the correct object due to limited relational perception.
@@
Many descriptions, such as the example in Fig. 1, involve multiple entities and relationships. Most existing methods perform relational perception only from pair-wise entities. The multi-modal fusion utilizes all relational information that corresponds with the local description (both “plant_9 and desk_43” and “plant_62 and desk_44” fulfill “the plant tucked in the furthest corner of the desk”). This may increase the uncertainty of the object representation. In cases where multiple entity pairs simultaneously adhere to the described relationships, local reasoning is insufficient for precise localization.
@@
Although graph-based models can represent both local neighborhoods and global contexts [chen2024survey], the scene graphs in current research establish connections between all spatially adjacent or semantically related entities. This also produces incorrect guidance in cross-modal understanding due to the aggregation of excessive noisy information in graph nodes. Our goal is to infer the n-ary combinations that simultaneously satisfy all referential conditions detailed in the description. This process involves a close relationship between the multi-modal fusion feature and n-ary global representation, while ensuring the removal of interference from noisy information.
@@ II Related Work
The research on 3D object grounding aims to localize the specific object using referential descriptions in 3D scenes. The current mainstream methods rely on the two benchmarks ScanRefer [chen2020scanrefer] and ReferIt3D [achlioptas2020referit3d] for further research. From a modeling structure perspective, these methods can be classified as two-stage and one-stage [liu2024survey]. The two-stage method first employs a 3D detector to identify object proposals and then utilizes a vision-language perception model to determine the target [chang2024mikasa, huang2022multi, zhang2023multi3drefer]. The one-stage method processes scene point clouds end-to-end and directly outputs the target position [wu2023eda, luo20223d, cai20223djcg]. Describing complex scenes often requires multiple spatial relationships to clarify object positions, making the accurate understanding of these relationships a primary issue. Zhao et al. [zhao20213dvg] and Chang et al. [chang2024mikasa] integrate pre-extracted spatial relationship matrices into the transformer to improve relation awareness. Cai et al. [cai20223djcg] improve geometric relationship expressions through the proposed attribute-relation enhancement module. Huang et al. [huang2021text] construct an instance graph to aggregate neighboring relationship information. Chen et al. [chen2022language] point out that modeling only neighboring objects in the graph ignores the distance expression of relationships and propose a spatial self-attention layer. The state-of-the-art methods [zhang2024cross, xu2024multi] focus more on better visual representations of objects from attributes or semantic labels. Additionally, some works utilize large language models for text analysis or data augmentation to enhance the language comprehension capabilities of models [guo2023viewrefer, chen2023unit3d, jia2024sceneverse, zhan2024mono3dvg]. However, these methods still struggle to make accurate distinctions when faced with multiple objects identical to the target category. In this paper, we explicitly model binary and n-ary relations to facilitate global multi-modal alignment, which enables more accurate localization of objects under complex circumstances.
@@ II-B Scene Graph in Vision-Language
Scene graphs provide a high-level understanding of visual scenes, typically representing objects as nodes and relational representations as edges [chen2024survey]. Because of the topological homogeneity of visual and textual modalities in scene graphs, scene graphs are widely employed in vision-language tasks to facilitate cross-modal alignment and joint understanding [chen2022multi, deng2023transvg++, yang2020relationship]. Graph neural networks (GNNs) can automatically learn the relationships between nodes when paired with scene graph structures. The introduction of the graph attention network (GAT) [velivckovic2017graph], which incorporates attention attributes, helps understand complex scenes [tu2023relation, chen2025multi, zhang2024question]. In the field of 3D visual understanding, researchers also construct scene graphs to enhance model understanding of geometric relationships [dhamo2021graph, armeni20193d, yan2023comprehensive]. Koch et al. [koch2024open3dsg] use the objects detected by the 3D detector as graph nodes and construct scene graphs with LLM filtering object-pair relationships as edges. Chen et al. [chen2024clip] predict the relationship score by concatenating the features of two objects and inputting them into a fully connected network, which is used to construct edges. Strader et al. [strader2024indoor] utilize logical rules in spatial ontology as prior knowledge, employing LLM to extract common spatial concepts and establish a scene graph. However, the above methods only focus on the spatial or semantic relationships between pairwise objects when constructing scene graphs. They cannot effectively adapt to the 3D localization task with multiple referential relationships. Our method predicts n-ary relationships globally to establish graph edges and enhance vision-language alignment.
@@ III-A Overview
The overall structure of our method is shown in Fig. 2. Descriptions and objects within 3D scenes are encoded into text and object features, respectively. Subsequently, the binary-to-n-ary progressive relational learning module (B2N-PRL) module anticipates entity relationships via a text-guided binary to n-ary modelling process. During the training phase, we employ the LLM to extract entity combinations from descriptions related to the referential dependencies, which are used to supervise the progressive relational learning. Most referential relationships involve no more than four entities. Therefore, our relational reasoning iterates only twice, with each n-ary combination comprising 2 to 4 objects. The top $K_{2}$ n-ary combinations with the highest responses are used to construct the scene graphs. The attention-driven graph learning module updates node features and localizes the target by confidence from a multilayer perceptron (MLP).
@@ III-B Object-Level Representation
Since entities are the basic units of relational learning, we prioritize object-level visual feature encoding within scene contexts. Following the two-stage detect-then-match approaches, the scene point cloud is decoded into instance-level point clouds via segmentation networks, then extracts object features using a 3D point cloud encoder. Inspired by [xiao2025as3d], our method employs a training-free segmentation model based on multiple views to generate class-agnostic object proposals. Simultaneously, we harness projection data during the segmentation process to derive 2D features from the pre-trained multi-modal model. The incorporation of pre-trained 2D visual features addresses the inadequate representation of surface color and texture details in point cloud encoding. This enables the model to effectively assimilate comprehensive multi-modal attributes within the constraints of limited 3D data. Therefore, we fuse the 2D features with 3D object features to enhance object-level visual representation as follows:
@@
where $F_{obj\_2D}$ and $F_{obj\_3D}$ represent the 2D object features from the pre-trained model CLIP [radford2021learning] and the 3D object features from the point cloud encoder, respectively. $F_{obj\_2D}$ is mapped by linear function $\phi$ to obtain fine-tuned features that have the same dimensions as $F_{obj\_3D}$. The MLP-based fusion of point cloud features with multi-modal 2D features may potentially obscure certain geometric information. To mitigate this issue, the enhanced features are subsequently integrated with the original point cloud features.
@@ III-C Progressive Relational Learning
The textual descriptions for the target with distractors often involve multiple referential relationships, which are also seen as the combination of several pairwise object relationships. However, existing models predominantly employ pairwise relationship modeling followed by direct fusion with text modalities, exhibiting insufficient global contextual awareness in descriptions with n-ary relational dependencies. We reformulate the object grounding task as a multi-relational structure inference problem. The target entity is localized by identifying candidate object combinations that satisfy the composite relational constraints. We utilize the LLM to extract binary relationships with textual representation about entity names. The corresponding objects in the scene are identified using these entity names to supervise the relational learning model. We design the progressive relational learning module, B2N-PRL, for reasoning about binary to n-ary relationships, as shown in Fig. 3.
@@ III-C2 Binary Relational Modeling
Binary relations as fundamental units for modeling complex spatial relationships are encoded from paired objects. We employ a BERT-based [kenton2019bert] textual encoder to process natural language sentences and obtain text features $T$. During the perception process, sentence decomposition is unnecessary, and the holistic textual representation T directly engages in multi-modal interactions with visual information at various stages. Similarly to previous work [chang2024mikasa], we improve the spatial perception ability of the model by integrating box embedding $F_{box}$ and artificially extracted geometric features $F_{geo}$. The specific calculation is as follows:
$\mathbf{O^{\prime}}_{i}=\mathbf{Att}_{cross}(\mathbf{O}_{i}+F_{box_{i}},\mathbf{T})$ (2)
$\mathbf{B}_{i,j}=\mathbf{O^{\prime}}_{i}\times\mathbf{O^{\prime}}_{j}+F_{geo_{i}j}$ (3)
@@
$L_{br}=-\frac{1}{N\times N}\sum_{i=1}^{N\times N}\left[r_{i}\log(S_{1,i})+(1-r_{i})\log(1-(S_{1,i})\right]$ (4)
where $r_{i}\in\{0,1\}$ indicates whether the $i$-th entity combination matches the description.
@@ III-D2 Graph-Structured Multi-Modal Network
We integrate 3D vision with text in a graph structure and propose a multi-modal attention network. As shown in Fig. 4, we employ cross-attention to facilitate interaction between object features and text features, and graph attention to aggregate node information concerning referential relationships through n-ary connections. Specifically, node features are fused with binary features obtained from the relation learning module using a self-attention mechanism to capture object-to-object dependencies. Then we utilize the multi-head cross-attention operation to fuse the textual information and achieve global perception within a graph attention layer. The graph attention operation updates the node values as follows:
@@ III-E Loss Function
During training, we set up the localization loss $L_{ref}$ to supervise the final target prediction results and use the classification loss $L_{t}$ and $L_{v}$ to guide the textual and visual encoding as in previous work [chen2022multi, guo2023viewrefer, chang2024mikasa, zhang2024cross]. It is noted that previous methods use original instance classes as the gt labels for $L_{v}$. We reduce the number of mapped classes to rely less on the classification information inherent in the visual modality. In addition to them, we propose binary and n-ary relational losses ($L_{br}$ and $L_{nr}$ in Section III-C) to supervise the B2N-PRL module. The whole loss function is grouped as:
@@ IV-B Implementation Details
We use GPT-4o mini [achiam2023gpt] as the LLM-based agent to extract entity combinations related to spatial relationships from descriptions. All multi-head cross-attention, self-attention, and graph attention layers in our methods are configured with eight heads. The loss weight coefficients $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ are set to 0.1, 0.5, and 2.0, respectively. The models are trained with a batch size of 20 for 150 epochs, using the Adam [kingma2014adam] optimizer. The learning rate is initialized to $5\times 10^{-4}$ with a decay of 0.65 every 10 epochs from 30 to 80. In the B2N-PRL module, the number of selected binary and n-ary relationships is set to $K_{1}$=16 and $K_{2}$=16.
@@
ReferIt3D [achlioptas2020referit3d] 2020 35.6 27.9 43.6 32.5 37.1 40.8 31.5 44.7 39.2 40.8
TGNN [huang2021text] 2021 37.3 30.6 44.2 35.8 38.0 45.0 36.9 48.5 45.8 45.0
3DVG-Trans [zhao20213dvg] 2021 40.8 34.8 48.5 34.8 43.7 51.4 44.9 54.2 44.6 51.7
FFL-3DOG [feng2021free] 2021 41.7 35.0 48.2 37.1 44.7 - - - - -
LAR [bakr2022look] 2022 48.9 42.3 58.4 47.4 52.1 59.4 51.2 63.0 50.0 59.1
SAT [yang2021sat] 2021 49.2 42.4 56.3 46.9 50.4 57.9 50.0 61.2 49.2 58.3
M3DRef-CLIP [zhang2023multi3drefer] 2023 49.4 43.4 55.6 42.3 52.9 - - - - -
3D-SPS [luo20223d] 2022 51.5 45.1 58.1 48.0 53.2 62.6 65.4 56.2 49.2 63.2
EDA [wu2023eda] 2023 52.1 - - - - 68.1 - - - -
BUTD-DETR [jain2022bottom] 2022 54.6 48.4 60.7 46.0 58.0 67.0 63.2 68.6 53.0 67.6
MVT [huang2022multi] 2022 55.1 49.1 61.3 54.3 55.4 64.5 58.8 66.9 58.4 64.7
ViewRefer [guo2023viewrefer] 2023 56.0 49.7 63.0 55.1 56.8 67.0 62.1 68.9 52.2 67.7
ViL3DRel [chen2022language] 2022 64.4 57.4 70.2 62.0 64.5 72.8 67.9 74.9 63.8 73.2
MiKASA [chang2024mikasa] 2024 64.4 59.4 69.7 65.4 64.0 75.2 67.3 78.6 70.4 75.4
MA2TransVG [xu2024multi] 2024 65.2 57.6 71.1 62.5 65.4 73.9 69.3 76.0 64.5 73.8
xM_Match [zhang2024cross] 2024 66.2 59.9 72.8 63.8 67.5 74.6 71.3 75.9 65.0 74.7
@@
3D-SPS [luo20223d] one-stage Pixel-N 29.8 37.0
BUTD-DETR [jain2022bottom] one-stage - 32.8 37.1
3DVLP [zhang2024vision] one-stage Pixel-N 33.4 40.5
EDA [wu2023eda] one-stage - 37.6 42.3
MCLN [qian2024multi] one-stage - 38.4 42.6
TGNN [huang2021text] two-stage 23.2 29.7
InstanceRefer [yuan2021instancerefer] two-stage - 24.8 32.9
ViewRefer [guo2023viewrefer] two-stage - 26.5 33.7
ViL3DRel [chen2022language] two-stage - 30.7 37.7
3D-LP [jin2023context] two-stage 32.2 38.1
xM_Match [zhang2024cross] two-stage - - 39.3
@@
ScanRefer [chen2020scanrefer] two-stage Pixel-N 21.1 27.4
3DVG-Trans [zhao20213dvg] two-stage Pixel-N 28.4 34.7
3DJCG [cai20223djcg] two-stage Pixel-N 30.8 37.3
D3Net [chen2022d] two-stage Pixel-N 30.1 37.9
3D-LP [jin2023context] two-stage Pixel-N 33.4 39.5
MA2TransVG [xu2024multi] two-stage Pixel-N 41.4 45.7
M3DRef-CLIP [zhang2023multi3drefer] two-stage Object-3 36.8 44.7